OpenStudy (anonymous):

Its a Polynomial question Multiply 9x(4x-13)

OpenStudy (anonymous):

we have : 9x * 4x +9x * (-13) = 36x^2+9x-13 Got it ? @shakur

OpenStudy (anonymous):

No no ! We have : -13 * 9x = -117x

OpenStudy (anonymous):

we have : 36x^2-117x

OpenStudy (anonymous):

I got 36x^2-117x

OpenStudy (anonymous):

9x * 4x +9x * (-13)

OpenStudy (anonymous):

@shakur

OpenStudy (anonymous):

Thanks Guyz ,But I have another polynomial question which I cant figure it the solution can you guys work it out for me

OpenStudy (anonymous):

Find the missing monomial -6x^10.?=84x^23

OpenStudy (anonymous):

Sure !!!

OpenStudy (anonymous):

What should we find ??? x ?!!!!!!

OpenStudy (anonymous):

the missing monomial of this question -6x^10. ? =84x^23

OpenStudy (anonymous):

K ! So divided 84 on 10 and x^23 on -6x^10 Have : 84x^23/-6x^10 = -14x^13 Got it ?:)